OpenStudy (anonymous):

6/36 in simpilest form ?

myininaya (myininaya):

what factors does 6 and 36 have in common?

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

zbay - hmmm.....did you not see myininaya was trying to help? Before simply providing your answer?

myininaya (myininaya):

i don't think the person was gonna answer anyways

OpenStudy (anonymous):

probably true.

OpenStudy (anonymous):

I did but i also noticed that the person she was trying to help didn't respond yet and still hasn't responded. So i was trying to save some problems for people that just want the answers anyways.

myininaya (myininaya):

Its all cool. We are all totally super happy! :)

OpenStudy (anonymous):

i actually dont know what the common factors are..i am absoulutley terrible at math and im failing thats why id didn't answer and because i wsnt paying attention i learn slower than others and my teachers are racisit so there not helping at all or trying to help me learn better

myininaya (myininaya):

Too much information about your teachers. 6 and 36 both have the common factor 6 since 6=6(1) and 36=6(6) 6/36=[6(1)]/[6(6)] a 6 cancels on top and bottom so we are left we 1/6

OpenStudy (anonymous):

thank you

myininaya (myininaya):

np

myininaya (myininaya):

did it make sense?

myininaya (myininaya):

do you think you can do 5/15 ?

myininaya (myininaya):

reduce 5/15

OpenStudy (anonymous):

yeah it made sense and i dont know ill try

myininaya (myininaya):

ok cool

OpenStudy (anonymous):

well 3 * 5 = 15 and 5*1 = 5 so 5 cancls out i dont know what else thou but good so far ?

myininaya (myininaya):

thats awesome so far

myininaya (myininaya):

so we have 5/15=[5(1)]/[5(3)]

myininaya (myininaya):

and you said the 5's cancel so that means we have what?

OpenStudy (anonymous):

4/5 you add the 1+3 ?

myininaya (myininaya):

after canceling the 5's you should be with left with 1 on top and 3 on bottom 1/3

myininaya (myininaya):

you put more work into it than it was nice try

myininaya (myininaya):

do you want to try another one?

myininaya (myininaya):

well if you decide you want to? try reducing 4/14